OpenStudy (anonymous):

How to show this "identity" for multiplication of series? \[\sum_{k=0}^{\infty}\sum_{n=-k}^{\infty}b_{k,n} = \sum_{n=-\infty}^{\infty}\sum_{k=\max{0,-n}}^{\infty}b_{k,n}\]

OpenStudy (kainui):

I'd probably play around with like finitely many of the terms and see what that looks like and then try to generalize it to an infinite case. At least I would convince myself that it was probably true even if I couldn't rigorously prove it.

OpenStudy (anonymous):

I was assuming there'd be a way to substitute something, but I would think that'd change the \(b_{k,n}\) part.. But I'm not really sure how you would get a sum to go from 0 to \(\infty\) to \(- \infty \) to \(\infty\). I figure that's what the -k part is for, but yeah.

OpenStudy (anonymous):

Im not really even sure how I'd write this out, to be honest. Should I be picking finite numbers for the indeces and then writing that out? I guess having that \(-\infty\) part makes me unsure how to write it.

OpenStudy (kainui):

Yeah beats me I was sorta distracted earlier but I'll try to see. My first attempt at visualization is to sorta sketch out a graph of the discrete points that we're summing over with coordinates (k,n) just to sorta have something to sorta picture in my mind.

OpenStudy (kainui):